OpenStudy (ihateschoolx.x):

(cos Θ + cos Θ)2 + (cos Θ + cos Θ)2 Please help me solve. Answer choices: sin2 Θ 4 8cos2Θ cos2 Θ

OpenStudy (radar):

Could you also write \[(2Cos \theta)^{2}+(2Cos \theta)^{2}=4Cos ^{2}\theta +4Cos ^{2} \] and take it on from there?

OpenStudy (ihateschoolx.x):

I don't know what to do from there

OpenStudy (ihateschoolx.x):

Is the answer 4?

OpenStudy (radar):

Ooop left off theta on last term. Could you just simply add and get \[8Cos ^{2}\theta \]

OpenStudy (radar):

Or am I just looking at to simply?

OpenStudy (ihateschoolx.x):

Who knows. You're probably right. I'll let you know in a few :)

OpenStudy (radar):

O.K. Good luck with your study, but luck alone is not enough......did you follow and agree with that method??

OpenStudy (ihateschoolx.x):

Yes and you were right. Thank you so much!